OpenStudy (anonymous):

c=j+10 to find the value of c when j=2. c=

OpenStudy (whovianchick):

c=j+10 c=2+10 c=12

OpenStudy (kaylardz0405):

thats right

OpenStudy (anonymous):

thx @whovianchick